# FluoCLIP: Stain-Aware Focus Quality Assessment in Fluorescence Microscopy

### v1

@@ Abstract
Accurate focus quality assessment (FQA) in fluorescence microscopy remains challenging, as the stain-dependent optical properties of fluorescent dyes cause abrupt and heterogeneous focus shifts. However, existing datasets and models overlook this variability, treating focus quality as a stain-agnostic problem. In this work, we formulate the task of stain-aware FQA, emphasizing that focus behavior in fluorescence microscopy must be modeled as a function of staining characteristics. Through quantitative analysis of existing datasets (FocusPath, BBBC006) and our newly curated FluoMix, we demonstrate that focus–rank relationships vary substantially across stains, underscoring the need for stain-aware modeling in fluorescence microscopy. To support this new formulation, we propose FluoMix, the first dataset for stain-aware FQA that encompasses multiple tissues, fluorescent stains, and focus variations. Building on this dataset, we propose FluoCLIP, a two-stage vision-language framework that leverages CLIP’s alignment capability to interpret focus quality in the context of biological staining. In the stain-grounding phase, FluoCLIP learns general stain representations by aligning textual stain tokens with visual features, while in the stain-guided ranking phase, it optimizes stain-specific rank prompts for ordinal focus prediction. Together, our formulation, dataset, and framework establish the first foundation for stain-aware FQA, and FluoCLIP achieves strong generalization across diverse fluorescence microscopy conditions.
@@ 1 Introduction
Focus degradation varies fundamentally differently between the two modalities. In bright-field microscopy, blur tends to be spatially uniform, allowing low-level cues such as edges or gradients to reliably indicate focus. In contrast, fluorescence microscopy exhibits stain-dependent and non-uniform blur, since each fluorophore has distinct emission, signal-to-noise ratio, and background fluorescence. The perceived sharpness therefore varies with stain type and optical conditions, violating the underlying assumptions of conventional focus metrics. This modality gap poses critical challenges for FQA models. Thus, simple edge-based models such as FocusLiteNN [19], which performs reliably on bright-field datasets like FocusPath [11], fail to generalize to fluorescence datasets such as BBBC006 [14] and our newly curated FluoMix (Fig. 1). Their reliance on spatial gradients ignores stain-dependent variations, leading to unstable performance across fluorophores and tissue types. These observations highlights the need for stain-aware modeling capable of handling the heterogeneous focus degradation unique to fluorescence imaging.
@@
Despite this need, existing datasets fail to capture the stain-dependent variability characteristic of fluorescence microscopy. FocusPath dataset [11], designed for bright-field imaging, assumes uniform optical conditions and lacks stain variation observed in fluorescence microscopy. BBBC006 dataset [14] contains homogeneous cell-line images acquired under tightly controlled in-vitro conditions where signal variation is minimal, offering limited biological and optical diversity. In contrast, tissue-level fluorescence microscopy involves complex three-dimensional structures, heterogeneous cell populations, and spatially varying stain distributions. These factors produce intricate light–matter interactions and non-uniform defocus patterns that reflect real challenges in biomedical imaging.
@@
To comprehensively capture these factors, we construct FluoMix, a new dataset for Stain-Aware FQA in flourescence microscopy. FluoMix contains multi-stain, tissue-level fluorescence images acquired under varying optical conditions, providing controlled yet diverse focus variations. By reflecting the spatial and biological heterogeneity of real tissue specimens, it establishes a practical foundation for developing and evaluating models capable of robust, stain-aware focus assessment.
@@
Building upon this dataset, we introduce FluoCLIP, a vision-language framework that extends CLIP-style multi-modal learning [16] to focus quality assessment in fluorescence microscopy. Unlike bright-field imaging, focus perception in fluorescence microscopy depends not only on spatial sharpness but also on the spectral and semantic properties of stains. This motivates the use of a multi-modal approach, where textual stain descriptions provide complementary cues that visual features alone cannot capture. However, pretrained CLIP encoders lack grounding in such domain-specific concepts. Textual stain descriptions (e.g., “DAPI” or “Alexa-488”) have no learned visual or semantic correspondence in the original CLIP vocabulary. As a result, the model cannot directly associate stain identity with the focus patterns characteristic of each fluorophore. Moreover, stain identity and focus quality are inherently entangled in fluorescence images, making it difficult for a single-stage alignment strategy to separate stain semantics from focus-related variations.
@@
To address these limitations, FluoCLIP adopts a two-stage learning strategy. In the stain-grounding phase, the model aligns learnable stain tokens with CLIP visual features, allowing the text-side representations to acquire fluorophore-specific semantics that are absent from the pretrained vocabulary. In the subsequent stain-guided ranking phase, the learned stain embeddings guide ordinal focus predictor, enabling the model to modulate its focus perception according to stain-dependent appearance changes. This progressive design mitigates the semantic mismatch inherent in the original CLIP text encoder and disentangles stain semantics from focus-related variations.
@@ 2.1 Microscopy Focus Quality Assessment
Early approaches to focus quality assessment (FQA) in microscopy adapted traditional image quality metrics such as entropy, histogram analysis, and power spectrum statistics. For example, Bray et al. [3] proposed the Power Log-Log Slope (PLLS) method, which analyzes the frequency spectrum to detect focus loss, while Koho et al. [12] utilized statistical measures to quantify sharpness. However, these hand-crafted methods are sensitive to staining variability and lack generalizability across imaging protocols.
With the rise of deep learning, data-driven methods have improved FQA robustness. Yang et al. [20] synthesized defocus using the point spread function and trained models with ordinal regression loss to reflect the ordered nature of focus levels. Albuquerque et al. [2] introduced a microscopy-specific ordinal loss, later applied to other biomedical tasks like cell development stage classification. Zhang et al. [21] further incorporated gradient maps and label distribution learning, while Gu [9] proposed training with sharp–blurred image pairs to enhance generalization.
Nonetheless, most of these models are trained on brightfield images with stable staining and gradual focus variation. In contrast, fluorescence microscopy presents abrupt focus shifts, light scattering, and stain-specific emission properties that complicate focus modeling. These challenges demand methods that can integrate both structural sharpness and stain semantics to improve focus quality estimation. This motivates exploring vision–language pretraining for microscopy focus modeling, where textual stain descriptions can complement purely visual FQA features.
2.2 CLIP for Ordinal Regression
Recent research has adapted CLIP [16] to ordinal regression by treating it as a vision-language matching problem. Each ordinal level is encoded as a prompt, and image classification is performed by measuring similarity in the joint embedding space.
OrdinalCLIP [13] introduced rank-aware prompting, where intermediate prompts are generated via linear interpolation between base rank prompts (e.g., “age 0” and “age 100”) in the text embedding space. This preserves ordinal structure but does not explicitly ensure semantic alignment between visual and textual modalities, limiting performance in tasks with fine-grained visual differences. To address this, L2RCLIP [18] proposed RankFormer, a token-level prompt tuning module, and introduced a cross-modal pairwise ranking loss to reinforce alignment between image features and ordinal prompt ordering. This model better captures semantic ordering across modalities. NumCLIP [6] further argued that CLIP lacks numerical understanding and proposed a two-stage learning scheme that disentangles numerical semantics from natural language, improving performance in score- or quantity-based prediction tasks.
Domain-specific variants such as BiomedCLIP [22] extend CLIP pretraining to biomedical datasets and improve alignment for radiology and pathology images. However, these models are trained on text–image pairs describing anatomical or disease entities, not stain-level fluorescence characteristics, and therefore remain limited in capturing stain semantics essential for focus modeling
While prior CLIP-based ordinal models have focused on natural or radiological imagery, our work extends this paradigm to fluorescence microscopy, where ordinal cues are intertwined with stain-dependent optical effects. To this end, we introduce a two-stage training framework that first learns stain-specific visual–textual alignment and then performs stain-conditioned rank learning, enabling robust and semantically consistent focus prediction across diverse fluorescence microscopy conditions.
@@ 3.1 Problem Statement
While this formulation enables CLIP to perform ordinal prediction via prompt-based classification, it inherently assumes that both visual and textual embeddings are semantically aligned in a domain-agnostic manner. In fluorescence microscopy, however, focus quality is governed by stain-dependent optical phenomena, including light scattering, photobleaching, and fluorophore emission variability, that generic CLIP prompts neither describe nor align with. This motivates a stain-aware adaptation that grounds textual embeddings in fluorescence-specific semantics.
@@ 3.2.1 Stage 1: Stain-Grounding
Given a predefined set of stains, each stain description is tokenized and embedded using CLIP’s token embedding layer. Let denote a collection of learnable stain embeddings, where each represents represents the token-level embedding sequence for stain (: the number of tokens per stain, : embedding dimension). These embeddings are grounded to the CLIP visual encoder to encode fluorophore-specific characteristics.
@@
where . The alignment is optimized in a way of predicting the stain label using the cross-entropy and KL-divergence objectives described in Sec. 3.3. This process aligns each image feature with the correct stain embedding, encouraging the model to capture fluorophore-specific semantics in the shared vision–language space.
@@ 3.2.2 Stage 2: Stain-Guided Ranking
For an input image with stain label , the stain-conditioned rank embeddings are inserted into pseudo-sentence templates (Fig. 2):
@@
where denotes cosine similarity, and is the visual embedding. These similarities are trained to predict the focus label using the cross-entropy and KL-divergence objectives described in Sec. 3.3, with the ground truth focus label .
@@ 4.1 Motivation and Design Principles
Consequently, existing datasets do not reflect the complexity of real-world fluorescence imaging, where multiple tissues, stains, and acquisition protocols coexist. To fill this gap, we introduce FluoMix—multi-tissue, multi-stain dataset specifically designed for stain-aware FQA in fluorescence microscopy.
@@ 4.2 FluoMix Composition
FluoMix aggregates fluorescence microscopy images across multiple tissue types and staining conditions to capture diverse optical and biological characteristics. Specifically, it includes images from brain, lung, and liver tissues, with up to four distinct stains per sample. Each field of view (FOV) is acquired as a complete 32-plane z-stack, covering full range from sharp to severely blurred slices. In total, FluoMix contains images spanning unique stain-tissue combinations.
This design allows FluoMix to explicitly model stain-dependent focus degradation, providing a realistic dataset for testing Stain-Aware FQA methods. Alongside the raw images, we provide detailed metadata describing the stain identity, target protein, and acquisition protocol, enabling multimodal research, including vision–language modeling, beyond conventional image-only baselines. Representative visual examples are provided in the Supplementary Material.
@@ 4.3 Annotation Protocol
We adopt a relative annotation scheme inspired by prior work [11]. For each z-stack, the visually best-focused slice is selected by a expert as a reference, and all other slices are assigned relative focus levels from (best focus) to (worst). In this release, we provide annotations with discrete ordinal levels obtained by uniformly partitioning the 32 slices.
This strategy offers a consistent yet scalable protocol for large-scale datasets, while maintaining ordinal relationships essential for FQA. Importantly, the full raw z-stacks are released so that future work can explore alternative anootation schemes-such as continuous scores, perceptual metrics, or data condensation approaches-without re-acquisition overhead. By decoupling dataset release from a fixed labeling protocol, FluoMix provides a flexible foundation for future research in stain-aware FQA modeling.
5 Empirical Analysis of Stain-Dependent Focus Behavior
Having introduced the dataset design, we next analyze how focus–rank relationships vary across FocusPath [11], BBBC006 [14], and our proposed FluoMix to empirically validate the need for stain-aware focus modeling. The spatial frequency (SF) metric quantifies the density of high-frequency components in an image. Since well-focused images preserve high-frequency details while defocus attenuates them, SF serves as an effective proxy for sharpness and focus quality. We therefore use the spatial frequency (SF) metric as a quantitative measure to compare how focus behavior differs between bright-field and fluorescence modalities. Further explanation about SF metric are provided in the Supplementary Material.
@@ 5.1 Per-Rank Behavior of Sharpness Metric
We first examine the relationship between the spatial frequency (SF) metric and the discrete focus ranks to verify that the metric reliably reflects the degree of defocus represented by the ranks. As shown in Fig. 33(a), the mean SF values exhibit a generally monotonic decrease as the focus level moves from in-focus to defocused regions across all three datasets. This confirms that the SF metric can reasonably capture the overall directionality of focus degradation implied by the ordinal labels, validating its use as a quantitative sharpness indicator for subsequent analysis. In other words, while focus ranks are discrete and manually defined, the SF metric provides as an objective, continuous indicator of focus quality, exhibiting a strong correlation with the discrete rank progression.
@@ 5.2 Stain-Wise Variability Across Datasets
We then analyze the distribution of the SF metric across different stains within each dataset in Fig. 3. In FocusPath [11], which consists of bright-field H&E-stained whole-slide images, the per-stain variance is modest, with stains exhibiting closely aligned SF ranges and largely overlapping distributions, indicating stain-invariant focus behavior. In contrast, BBBC006 [14] and FluoMix exhibit substantial stain-dependent variability. Different fluorescent stains occupy distinct SF ranges with largely non-overlapping variances, demonstrating that focus degradation is highly stain-dependent in fluorescence microscopy. This observation supports the hypothesis that focus quality should be modeled as a function of stain characteristics rather than assuming a universal focus-rank mapping.
@@ 5.3 Sharpness Metric and Rank Correlation
To further quantify this effect, we compute the Spearman rank correlation coefficient (SRCC) between focus rank and SF metric for each stain. Since spatial frequency decreases as defocus increases, SF monotonically decreases with rank. This produces an inherently negative SRCC value, where the absolute magnitude reflects the strength of the relationship: larger values indicate that SF aligns more consistently with the underlying focus levels. As shown in Table. 1, FocusPath [11] achieves the highest correlation, indicating that its ordinal labels correspond closely to actual blur magnitudes under stain-invariant conditions. However, correlations drop significantly in BBBC006 [14] and FluoMix, revealing that rank–blur consistency weakens in fluorescence microscopy due to heterogeneous optical properties across stains. This reduced correlation implies that a single, stain-agnostic ranking model cannot reliably capture focus variation across fluorescence conditions.
@@
Summary. These analyses demonstrate that while bright-field microscopy exhibits uniform, stain-independent focus behavior, fluorescence microscopy shows strong stain-specific trends and inconsistent rank–focus correlations. Therefore, accurate focus quality assessment in fluorescence microscopy requires a stain-aware ordinal framework capable of modeling these stain-conditioned focus variations.

### v2

@@ Abstract
Accurate focus quality assessment (FQA) in fluorescence microscopy is challenging due to stain-dependent optical variations that induce heterogeneous focus behavior across images. Existing methods, however, treat focus quality as a stain-agnostic problem, assuming a shared global ordering. We formulate stain-aware FQA for fluorescence microscopy, showing that focus-rank relationships vary substantially across stains due to stain-dependent imaging characteristics and invalidate this assumption. To support this formulation, we introduce FluoMix, the first dataset for stain-aware FQA spanning multiple tissues, fluorescent stains, and focus levels. We further propose FluoCLIP, a two-stage vision-language framework that grounds stain semantics and enables stain-conditioned ordinal reasoning for focus prediction, effectively decoupling stain representation from ordinal structure. By explicitly modeling stain-dependent focus behavior, FluoCLIP consistently outperforms both conventional FQA methods and recent vision-language baselines, demonstrating strong generalization across diverse fluorescence microscopy conditions. Code and dataset are publicly available at https://fluoclip.github.io/.
@@ 1 Introduction
Focus degradation varies fundamentally between the two modalities. While bright-field microscopy blur is often spatially uniform, fluorescence microscopy exhibits stain-dependent and non-uniform blur due to differences in emission spectra, signal-to-noise ratio, and background fluorescence. This behavior arises from the wavelength-dependent diffraction limit (, where denotes the numerical aperture), leading to resolution differences across fluorophores with different emission wavelengths (). In our FluoMix (), fluorophores ranging from 405 nm (“DAPI”) to 647 nm (“Alexa 647”) exhibit a 1.6 variation in theoretical resolution. As a result, even perfectly focused images (Rank 0) inherently possess different high-frequency characteristics across stains, violating the underlying assumptions of stain-agnostic focus metrics.
@@
This modality gap poses critical challenges for FQA models. Edge-based approaches such as FocusLiteNN [22], which performs well on bright-field datasets like FocusPath [13], fail to generalize to fluorescence datasets such as BBBC006 [17] and FluoMix (Figure 1). Their reliance on spatial gradients ignores stain-specific optical regimes, leading to unstable performance. These observations highlight the need for stain-aware modeling capable of handling heterogeneous focus degradation in fluorescence imaging.
@@
Despite this need, existing datasets fail to capture the stain-dependent variability characteristic of fluorescence microscopy. FocusPath [13], designed for bright-field imaging, assumes uniform optical conditions and lacks stain variation observed in fluorescence microscopy. BBBC006 [17] contains homogeneous cell-line images acquired under tightly controlled in-vitro conditions where signal variation is minimal, offering limited biological and optical diversity. In contrast, tissue-level fluorescence microscopy involves complex three-dimensional structures, heterogeneous cell populations, and spatially varying stain distributions. These factors produce intricate light–matter interactions and non-uniform defocus patterns that reflect real challenges in biomedical imaging.
@@
To comprehensively capture these factors, we construct FluoMix, a new dataset for Stain-Aware FQA in fluorescence microscopy. FluoMix contains multi-stain, tissue-level fluorescence images acquired under varying optical conditions, providing controlled yet diverse focus variations. By reflecting the spatial and biological heterogeneity of real tissue specimens, it establishes a practical foundation for developing and evaluating models capable of robust, stain-aware focus assessment.
@@
Building upon this dataset, we introduce FluoCLIP, a vision-language framework for fluorescence FQA. Unlike bright-field imaging, focus perception depends not only on spatial sharpness but also on stain-dependent appearance, motivating semantic conditioning via text. However, pretrained CLIP [19] lacks grounding in domain-specific stain concepts (e.g., “DAPI” or “Alexa-488”), and a single-stage alignment cannot disentangle stain semantics from focus-related variations.
@@
To address these limitations, FluoCLIP adopts a two-stage learning strategy. In the stain-grounding phase, learnable stain tokens are aligned with visual features to acquire fluorophore-specific semantics. In the subsequent stain-guided ranking phase, these learned stain embeddings guide ordinal focus prediction, enabling stain-conditioned focus modeling.
@@ 2.1 Microscopy Focus Quality Assessment
Early FQA methods adapted handcrafted metrics such as entropy or power spectrum statistics (e.g., PLLS [3], statistical sharpness measures [14]) to detect focus loss, but lack generalizability across imaging protocols. Recent deep learning approaches improve robustness via ordinal regression [23], specialized loss functions [2], label distribution learning [24], and sharp–blurred pair training [11]. However, these models are primarily optimized for bright-field imaging and remain stain-agnostic, failing to capture the stain-dependent focus behavior inherent in fluorescence microscopy. This gap necessitates modeling that integrates structural sharpness with stain-specific optical context.
2.2 CLIP-based Ordinal Regression
CLIP [19] has been increasingly adapted for ordinal tasks by treating rank prediction as a vision–language matching problem. Prior works preserve ordinal structure through rank-aware prompting or interpolation, such as OrdinalCLIP [15] and L2RCLIP [21], while NumCLIP [6] improves numerical reasoning in score-based prediction tasks.
Domain-specific variants such as BiomedCLIP [25], PubMedCLIP [8], and PMC-CLIP [16] are pretrained CLIP models tailored for biomedical domains rather than ordinal regression methods. While they improve alignment for biomedical imagery, they are trained on semantic image–text pairs and do not explicitly model stain-dependent optical variations or ordinal focus structure required for FQA.
@@ 3.1 Problem Statement
While this formulation enables CLIP to perform ordinal prediction, it assumes that both visual and textual embeddings are semantically aligned in a domain-agnostic manner. In fluorescence microscopy, however, focus quality is inherently stain-dependent due to wavelength dependent point spread function (PSF), leading to different degradation patterns across stains. As a result, a single global ordinal mapping becomes physically inconsistent. This motivates a stain-aware adaptation that grounds textual embeddings in fluorescence-specific semantics.
@@ 3.2.1 Stage 1: Stain-Grounding
Given a predefined set of stains, each stain description is tokenized and embedded using CLIP’s token embedding layer. Let denote a collection of learnable stain embeddings, where each represents the token-level embedding sequence for stain (: the number of tokens per stain, : embedding dimension). These embeddings are grounded to the CLIP visual encoder to encode fluorophore-specific characteristics.
@@
where . The alignment is optimized in a way of predicting the stain label using the cross-entropy and KL-divergence objectives described in Section 3.3. This process aligns each image feature with the correct stain embedding, encouraging the model to capture fluorophore-specific semantics in the shared vision–language space.
@@ 3.2.2 Stage 2: Stain-Guided Ranking
For an input image with stain label , the stain-conditioned rank embeddings are inserted into pseudo-sentence templates (Figure 2):
@@
where denotes cosine similarity, and is the visual embedding. These similarities are trained to predict the focus label using the cross-entropy and KL-divergence objectives described in Section 3.3, with the ground truth focus label .
@@ 4.1 Motivation and Design Principles
Consequently, existing datasets do not reflect the complexity of real-world fluorescence imaging, where multiple tissues, stains, and acquisition protocols coexist. To fill this gap, we introduce FluoMix—a multi-tissue, multi-stain dataset specifically designed for stain-aware FQA in fluorescence microscopy.
@@ 4.2 FluoMix Composition
FluoMix aggregates fluorescence microscopy images across multiple tissue types and staining conditions to capture diverse optical and biological characteristics. Specifically, it includes images from brain, lung, and liver tissues, with up to four distinct stains per sample. Each field of view (FOV) is acquired as a complete 34-plane z-stack, covering full range from sharp to severely blurred slices. In total, FluoMix contains approximately 56,000 images spanning 28 distinct imaging channels defined by stain–target–tissue combinations. These include 3 tissue types, 6 fluorophores, and 13 biological targets covering diverse functional categories such as neuronal, vascular, microglial, and epithelial markers. This diversity introduces heterogeneous optical regimes, making focus behavior inherently stain-dependent. This composition results in a broad range of optical and biological variability, reflecting diverse fluorescence imaging conditions (see Table 8 and Section 9.4).
This design allows FluoMix to explicitly model stain-dependent focus degradation, providing a realistic dataset for testing Stain-Aware FQA methods. Alongside the raw images, we provide detailed metadata describing the stain identity, target protein, and acquisition protocol, enabling multimodal research, including vision–language modeling, beyond conventional image-only baselines. Representative visual examples are provided in supplementary material.
@@ 4.3 Annotation Protocol
We adopt a relative annotation scheme inspired by prior work [13]. For each z-stack, the visually best-focused slice is selected by an expert as a reference, and all other slices are assigned relative focus levels from (best focus) to (worst). In this release, we provide annotations with discrete ordinal levels derived from the relative position within the z-stack.
This strategy offers a consistent yet scalable protocol for large-scale datasets, while maintaining ordinal relationships essential for FQA. Importantly, the full raw z-stacks are released so that future work can explore alternative annotation schemes such as continuous scores, perceptual metrics, or data condensation approaches-without re-acquisition overhead. By decoupling dataset release from a fixed labeling protocol, FluoMix provides a flexible foundation for future research in stain-aware FQA modeling.
5 Empirical Evidence for Stain-Dependent Focus Behavior
Having introduced the dataset design, we next analyze how focus–rank relationships vary across FocusPath [13], BBBC006 [17], and our proposed FluoMix to empirically validate the need for stain-aware focus modeling. The spatial frequency (SF) metric [7] quantifies the density of high-frequency components in an image. Since well-focused images preserve high-frequency details while defocus attenuates them, SF serves as an effective proxy for sharpness and focus quality. We therefore use the spatial frequency (SF) metric as a quantitative measure to compare how focus behavior differs between bright-field and fluorescence modalities. Since the point spread function (PSF) determines the spatial frequency response of an imaging system, wavelength-dependent PSF directly affects the high-frequency content captured by the SF metric. Further explanation of SF metric are provided in the supplementary material.
@@ 5.1 Per-Rank Behavior of Sharpness Metric
We first examine the relationship between the spatial frequency (SF) metric and the discrete focus ranks to verify that the metric reliably reflects the degree of defocus represented by the ranks. As shown in Figure 33(a), the mean SF values exhibit a generally monotonic decrease as the focus level moves from in-focus to defocused regions across all three datasets. This confirms that the SF metric can reasonably capture the overall directionality of focus degradation implied by the ordinal labels, validating its use as a quantitative sharpness indicator for subsequent analysis. In other words, while focus ranks are discrete and manually defined, the SF metric provides an objective, continuous indicator of focus quality, exhibiting a strong correlation with the discrete rank progression.
@@ 5.2 Stain-Wise Variability Across Datasets
We then analyze the distribution of the SF metric across different stains within each dataset in Figure 3. In FocusPath [13], which consists of bright-field H&E-stained whole-slide images, the per-stain variance is modest, with stains exhibiting closely aligned SF ranges and largely overlapping distributions, indicating stain-invariant focus behavior. In contrast, BBBC006 [17] and FluoMix exhibit substantial stain-dependent variability. Different fluorescent stains occupy distinct SF ranges with largely non-overlapping variances, demonstrating that focus degradation is highly stain-dependent in fluorescence microscopy. This observation supports the hypothesis that focus quality should be modeled as a function of stain characteristics rather than assuming a universal focus-rank mapping.
@@ 5.3 Sharpness Metric and Rank Correlation
To further quantify this effect, we compute the Spearman rank correlation coefficient (SRCC) between focus rank and SF metric for each stain. Since spatial frequency decreases as defocus increases, SF monotonically decreases with rank. This produces an inherently negative SRCC value, where the absolute magnitude reflects the strength of the relationship: larger values indicate that SF aligns more consistently with the underlying focus levels.
As shown in Table 1, FocusPath [13] achieves the highest correlation, indicating that its ordinal labels correspond closely to actual blur magnitudes under stain-invariant conditions. However, correlations drop significantly in BBBC006 [17] and FluoMix, revealing that rank–blur consistency weakens in fluorescence microscopy due to heterogeneous optical properties across stains. This reduced correlation implies that a single, stain-agnostic ranking model cannot reliably capture focus variation across fluorescence conditions.
This inconsistency arises because identical focus levels can produce different spatial frequency profiles across stains, causing a global ranking function to misinterpret stain-dependent sharpness variations. Fig. 4 provides visual evidence that images at the same ordinal focus level exhibit distinct appearance patterns across stains. These findings highlight the necessity of stain-aware modeling for reliable focus assessment in fluorescence microscopy.